OpenStudy (i_love_my_nieces):

What is the slope of the line that goes though (6, 5) and (3, 1)? 1. 3/4 2. -3/4 3. 4/3 4. -4/3

OpenStudy (anonymous):

You will want to use the slope formula which is y2-y1/x2-x1

OpenStudy (anonymous):

to get slope you do y2-y1/ x2-x1 .. you would do (5-1)/(6-3) which is 4/3

OpenStudy (anonymous):

So when plugged in your equation will read: 1-5/3-6

OpenStudy (i_love_my_nieces):

Ok

OpenStudy (mathstudent55):

The slope of a line calculated from tbhe coordinates of two points on the line is: \(\bf slope = \dfrac{\bf difference~~ in~~ y}{\bf difference~~in~~x} \)

OpenStudy (mathstudent55):

Calculate the difference in the y-values. Calculate the difference in the x-values. Divide the first difference by the second difference.

OpenStudy (mathstudent55):

You must make sure you calculate both differences the same way. That means, if the points are (a, b) and (c, d) and you decide to do the difference in y as d - b, then the difference in x must be in the same order, and it has to be c - a. If you want to do the difference in y as b - d, then the difference in x must be a - b.

OpenStudy (mathstudent55):

Let's do an example.

OpenStudy (i_love_my_nieces):

I think B is the answer.

OpenStudy (mathstudent55):

Slope of line through points: (4, 3) and (2, 6) diff in y = 6 - 3 = 3 diff in x = 2 - 4 = -2 slope = (diff in y)/(diff in x) = 3/(-2) = -3/2

OpenStudy (mathstudent55):

Now your problem. What is the difference in y? 5 - 1 = 4 What is the difference in x? 6 - 3 = 3 slope = (diff in y)/(diff in x) = 4/3 Which choice is that?

OpenStudy (i_love_my_nieces):

B

OpenStudy (i_love_my_nieces):

I mean C

OpenStudy (mathstudent55):

Isn't B -3/4?

OpenStudy (mathstudent55):

Oh, ok, that's better.

OpenStudy (mathstudent55):

Answer is C.